OpenStudy (jake124):

write 17/4 as a mixed number

OpenStudy (welshfella):

first work out 17 / 4 what is the quotient and what is the remainder?

OpenStudy (jake124):

4.25

OpenStudy (welshfella):

Ok I#m perhaps not explaining myself very well What is the whole number you get when you divide 17 by 4 ?

OpenStudy (jake124):

4

OpenStudy (welshfella):

right and the remainder is 17-4(4) = 17 - 16 = 1

OpenStudy (welshfella):

so the mixed number consists of the while number part which is 1 and the fraction is the reminder / 4 which = 1/4 so the mixed number is 4 1/4

OpenStudy (jake124):

thanks

OpenStudy (welshfella):

yw